# Graph-based Retrieval Augmented Generation for Dynamic Few-shot Text Classification

### v1

@@ Abstract.
Text classification is a fundamental task in natural language processing, pivotal to various applications such as query optimization, data integration, and schema matching. While neural network-based models, such as CNN and BERT, have demonstrated remarkable performance in text classification, their effectiveness heavily relies on abundant labeled training data. This dependency makes these models less effective in dynamic few-shot text classification, where labeled data is scarce, and target labels frequently evolve based on application needs. Recently, large language models (LLMs) have shown promise due to their extensive pretraining and contextual understanding. Current approaches provide LLMs with text inputs, candidate labels, and additional side information (e.g., descriptions) to predict text labels. However, their effectiveness is hindered by the increased input size and the noise introduced through side information processing. To address these limitations, we propose a graph-based online retrieval-augmented generation framework, namely GORAG, for dynamic few-shot text classification. GORAG constructs and maintains an adaptive information graph by extracting side information across all target texts, rather than treating each input independently. It employs a weighted edge mechanism to prioritize the importance and reliability of extracted information and dynamically retrieves relevant context using a minimum-cost spanning tree tailored for each text input. Empirical evaluations demonstrate that GORAG outperforms existing approaches by providing more comprehensive and accurate contextual information.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at ***.
@@ 1. Introduction
Text classification is a fundamental task that is connected to various other tasks, such as query optimization (Yu et al., 2023; Yu and Litchfield, 2020) data integration (Dong and Rekatsinas, 2018), and schema matching (Sahay et al., 2020; Zhang et al., 2024a). For example, in schema matching tasks, a text classifier can be trained to determine whether two schema attributes refer to the same concept based on their descriptions or metadata (Zhang et al., 2024a). In recent years, Neural Network (NN)-based models (Johnson and Zhang, 2015; Prieto et al., 2016; Liu et al., 2016; Dieng et al., 2017; Wang et al., 2017; Johnson and Zhang, 2017; Peters et al., 2018; Wang et al., 2018; Yang et al., 2019; Miao et al., 2021), such as CNN (Kim, 2014), Bert (Devlin, 2018) and RoBERTa (Liu, 2019), have demonstrated impressive performance on text classification tasks. However, the effectiveness of these NN-based approaches depend on abundant labeled training data, which requires significant time and human efforts (Meng et al., 2018). Consequently, these methods perform poorly with limited labeled text data (Xia et al., 2021), such as in few-shot text classification tasks (Wang et al., 2020; Aljehani et al., 2024).
Figure 1. An illustration of the structural composition of current Reteieval Augmented Generation based few-shot text classification models, where the few-shot labeled training data is indexed as an external information source. During classification, the model retrieves side information from the external information source to provide additional context that aids the classification process.
Also, in real-world applications, such as Web Of Science (Kowsari et al., 2017) and IMDb (imd, [n.d.]), target labels for text often change based on the application’s requirements (Xia et al., 2021), leading to dynamic text classification. For instance, movie labels might initially be {Action, Thriller, Comedy}, but platforms may need to add more genres (e.g., Romantic and Comic) to extend the original labels. Therefore, it is important to classify text into these new categories. Consequently, how to develop methods to dynamically classify text with limited labeled data (i.e., dynamic few-shot text classification) remains a necessity and open problem.
@@
Depending the technique for dynamic few-shot text classification task, current models can be categorized into three types, i.e., data augmentation-based (Meng et al., 2018; Meng et al., 2020; Xia et al., 2021), meta learning-based (Bansal et al., 2020; Zhang et al., 2023a), and large language model (LLM)-based approaches (Chen et al., 2023; Sarthi et al., 2024; Edge et al., 2024; Gutiérrez et al., 2024; Guo et al., 2024). Firstly, data augmentation-based models create additional training data by mixing the pairs of the few-shot labeled data and assign a mixed labels for these created data based on the labels of each pairs of data (Xia et al., 2021). Also, several data augmentation approaches create extra semantic related content based on the label names (Meng et al., 2018; Meng et al., 2020). However, due to the limited labeled data, the generated text data has very limited patterns. Consequently, the text classification models trained on these generated data are over-fitting on limited text data and are not generalizable (Liu et al., 2023). Also, meta-learning models (Bansal et al., 2020; Zhang et al., 2023a) train meta-learners on each base class to quickly adapt to new classes with minimal data. However, meta-learning still requires a substantial amount of labeled data for base classes.
@@
Recently, large language models (Sun et al., 2021; Jiang et al., 2023; Touvron et al., 2023; Achiam et al., 2023; Yang et al., 2024), pretrained on extensive corpora, have achieved significant success due to their superior and comprehensive text understanding abilities. Several researchers (Chen et al., 2023; Sarthi et al., 2024) provide the text, candidate labels, and retrieve side information (e.g., descriptions of text and labels, or documents) for LLMs, and these LLMs output the labels of the text. Current work (Gao et al., 2023) demonstrates that this side information can provide context for the text and labels, enabling LLMs to comprehensively understand these inputs, which is a key component of the success of LLMs in text classification. However, the incorporation of side information can further increase the input size and noise (Chen et al., 2023), which impedes the efficiency and effectiveness of LLMs (Sarthi et al., 2024). As a result, compression-based retrieval-augmented generation (RAG) approaches (Edge et al., 2024; Gutiérrez et al., 2024; Guo et al., 2024) are proposed to construct graphs by extracting key information from the side information sources. These models then only retrieve the information (e.g., graph path or subgraphs) relevant to the target text as the side information for LLMs, reducing the length of inputs for LLMs.
However, existing compression-based RAG approaches (Edge et al., 2024; Guo et al., 2024; Gutiérrez et al., 2024) still have three limitations in the dynamic few-shot text classification task. Firstly, these approaches only extract information and build an information graph based on the side information of each text individually. In this way, they do not merge the information graphs of different text inputs into a single graph to obtain more comprehensive information for incoming target texts. Consequently, they fail to consider the correlation between side information of different text inputs, providing insufficient and incomplete context for LLMs. Secondly, existing approaches construct graphs by indexing extracted information linkages uniformly. They do not consider the varying importance and extraction confidence of each link, which may provide incorrect and unreliable context for LLMs. Thirdly, existing approaches select relevant information for each input text based on a globally predefined threshold. However, the optimal retrieval threshold can vary across different text samples, making the globally predefined threshold suboptimal for the entire dataset.
@@
To address the aforementioned issues, we propose a novel Graph-based Online Retrieval Augmented Generation framework for dynamic few-shot text classification, called GORAG. In general, GORAG constructs and maintains an adaptive online information graph by extracting side information from all target texts, tailoring graph retrieval for each input. Firstly, GORAG extracts keywords from the text using the LLM and links these keywords with the text’s ground truth label to represent the relationship between keywords and labels. Then, GORAG employs an edge-weighting mechanism when indexing edges into the information graph, assigning edge weights based on the importance of the keywords and their relevance to the respective text’s label. Secondly, when retrieving information from the graph for each text, GORAG retrieves candidate labels by constructing a minimum-cost spanning tree of text keywords mapped to the graph. After retrieval, the candidate labels serve as a filtered subset of the original target labels, which are then used to create input for the LLM. Since the generated spanning tree is determined solely by the graph information and the extracted text keywords, GORAG achieves adaptive retrieval without relying on any human-defined retrieval thresholds.
We summarize the novel contributions as follows.
  • We present a RAG framework for few-shot text classification tasks, namely GORAG. The GORAG framework consists of four stages: (1) Graph construction, where a graph is built with keywords from the training texts considering their correlation with the texts’ label; (2) Candidate label retrieval, where relevant labels are adaptively retrieved from the graph for label pre-filtering; (3) Prompt construction, where prompts are generated based on the filtered candidate labels and the testing texts; and (4) Online indexing, where the keywords of the un-labeled text are indexed to the constructed graph based on the prediction result.
@@
  • We develop a novel graph edge weighting mechanism based on the keyword nodes’ importance within the text corpus, which can be applied during both graph construction and online indexing. This mechanism enables our approach to effectively model the relevance between keywords and labels, thereby reducing the noise introduced to the graph.
@@
  • To avoid the need for human-selected thresholds during retrieval, we formulate the candidate label retrieval problem which akin to the NP-hard Steiner Tree problem. To solve this problem, we modified the Mehlhorn algorithm (Mehlhorn, 1988) which provides an efficient and effective solution to the problem to generate candidate types from the information graph.
In the rest of this paper, we first present prelimnary and related work in section 2, the pipeline of our model GORAG in section 3, our experiments in section 4 and the final conclusion in section 5.
@@ 2. Preliminary and Related Works
In this section, we first introduce the preliminaries of dynamic few-shot text classification in Section 2.1 and then discuss the existing approaches in Section 2.2. The important notations used in this paper are listed in Table 1.
2.1. Dynamic Few-shot Text Classification
@@
Text classification (Kowsari et al., 2019; Minaee et al., 2021; Li et al., 2022) is a key task in real-world application that involves assigning predefined labels $y\in\mathcal{Y}$ to text $t=(w_{1},\cdots,w_{|t|})$ based on its words $w_{i}$. It is widely applied in areas like sentiment analysis (Kouadri et al., 2020; maamar2022sa),query optimization (Yu et al., 2023; Yu and Litchfield, 2020) data integration (Dong and Rekatsinas, 2018), and schema matching (Sahay et al., 2020; Zhang et al., 2024a). Traditional approaches rely on large labeled datasets (Devlin, 2018; Liu, 2019), which may not always be available. Recently, few-shot learning addresses this limitation by enabling models to classify text with only a small number of labeled examples per class (Meng et al., 2018; Meng et al., 2020). Dynamic classification (Parisi et al., 2019; Hanmo et al., 2024; Wang et al., 2024) introduces an additional challenge where new classes are introduced over multiple rounds, requiring the model to adapt to new classes while retaining knowledge of previously seen ones. Combining these aspects, Dynamic Few-Shot Text Classification (DFSTC) (Xia et al., 2021) allows the model to handle evolving classification tasks with minimal labeled data.
@@
In DFSTC, the model is provided with multiple rounds of new class updates. Specifically, In each round $r$, a new set of classes $\mathcal{Y}_{new}^{r}$ is introduced and the labelled dataset for $\mathcal{Y}_{\text{new}}^{r}$ is denoted as $\mathcal{D}_{new}^{r}=\cup_{y_{i}\in\mathcal{Y}_{\text{new}}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$, where per class $y\in\mathcal{Y}_{\text{new}}^{r}$ only has $k$ labeled examples $\{t_{j},y_{i}\}_{j=1}^{k}$. Also, we denote candidate cumulative labels from the first round to the $r$-th round as $\mathcal{Y}^{r}=\bigcup_{i=1}^{r}\mathcal{Y}_{\text{new}}^{i}$. Formally, at the $r$-th round, given the candidate labels $\mathcal{Y}^{r}$ and all labeled data $\mathcal{D}^{r}=\cup_{i=1}^{r}\mathcal{D}_{new}^{i}$, the target of DFSTC task is to learn a function $f^{r}_{\theta}$, which can learn scores for all target labels $\mathbf{s}^{r}_{k}={f}^{r}_{\theta}(t,\mathcal{Y}^{r})\in[0,1]^{|\mathcal{Y}^{r}|}$ for the unseen text $t_{k}$. Then, we can get the predicted label $y^{r*}_{k}\in\mathcal{Y}^{r}$ for the unseen text $t_{k}$ as follows.
(1) $y^{r*}_{k}=\arg\max_{y^{r}_{k}\in\mathcal{Y}^{r}}({f}_{\theta}(t,\mathcal{Y}^{r})[y^{r}_{k}])$
DFSTC is valuable for its ability to learn from limited labeled data, adapt to evolving class distributions, and address real-world scenarios where categories and data evolve over time. For example, in Round 1, the model might be trained on classes like Sports, Politics, and Technology. In Round 2, new classes such as Health and Education are introduced. The model must now classify texts into all five classes while retaining its knowledge of earlier ones. For instance, given the text ”The new vaccine shows promising results,” the model should classify it as Health.
2.2. Dynamic Few-shot Text Classification Models
Current dynamic few-shot text classification models can be broadly categorized into two types: Meta learning-based models and Data Augmentation-based models, and Large Language Model (LLM)-based models.
2.2.1. Meta learning-based models
Meta learning-based models train a meta-learner on base classes $C_{b}$ and then adapt it to new classes in each round. For instance, LEOPARD (Bansal et al., 2020) use the meta-learner to generate model parameters to predict classes in each round. ConEntail (Zhang et al., 2023a) trains the meta-learner to classify texts based on their nested entailments with class label names at each round. However, these models require a sufficient amount of training data to train the meta-learner, which differs from our experimental setting and limits their generalization ability.
2.2.2. Data Augmentation-based models

### v2

# GORAG: Graph-based Retrieval Augmented Generation for Dynamic Few-shot Text Classification

@@ Abstract.
Text classification is a fundamental task in data mining, pivotal to various applications such as tabular understanding and recommendation. Although neural network-based models, such as CNN and BERT, have demonstrated remarkable performance in text classification, their effectiveness heavily relies on abundant labeled training data. This dependency makes these models less effective in dynamic few-shot text classification, where labeled data is scarce, and new target labels frequently appear based on application needs. Recently, large language models (LLMs) have shown promise due to their extensive pretraining and contextual understanding ability. Current approaches provide LLMs with text inputs, candidate labels, and additional side information (e.g., descriptions) to classify texts. However, their effectiveness is hindered by the increased input size and the noise introduced through side information processing. To address these limitations, we propose a graph-based online retrieval-augmented generation framework, namely GORAG, for dynamic few-shot text classification. Rather than treating each input independently, GORAG constructs and maintains a weighted graph by extracting side information across all target texts. In this graph, text keywords and labels are represented as nodes, with edges indicating the correlations between them. To model these correlations, GORAG employs an edge weighting mechanism to prioritize the importance and reliability of extracted information and dynamically retrieves relevant context using a minimum-cost spanning tree tailored for each text input. Empirical evaluations demonstrate that GORAG outperforms existing approaches by providing more comprehensive and precise contextual information.
Large language models, online retrieval-augmented generation, few-shot learning
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/2025/02ccs: Information systems Clustering and classificationccs: Information systems Retrieval models and ranking
@@ 1. Introduction
Text classification is a fundamental task of text data mining and is connected to various tasks, such as tabular understanding (Chen et al., 2024c; Wang et al., 2024c; Wen et al., 2024a, b; He et al., 2025; Zheng et al., 2023), and recommendation (Du et al., 2024a; Han et al., 2024b; Wang et al., 2024b; Li et al., 2023; Zheng et al., 2024; Chiang et al., 2023; Du et al., 2024b). In recent years, Neural Network (NN)-based models (Johnson and Zhang, 2015; Prieto et al., 2016; Liu et al., 2016; Dieng et al., 2017; Wang et al., 2017; Johnson and Zhang, 2017; Peters et al., 2018; Wang et al., 2018; Yang et al., 2019; Miao et al., 2021), such as CNN (Kim, 2014), Bert (Devlin, 2018) and RoBERTa (Liu, 2019), have demonstrated impressive performance on text classification tasks. However, the effectiveness of these NN-based approaches depends on abundant labeled training data, which requires significant time and human efforts (Meng et al., 2018). Furthermore, in real-world applications, such as Web Of Science (Kowsari et al., 2017) and IMDb (imd, [n.d.]), target labels for text often change based on the application’s requirements (Xia et al., 2021), leading to the dynamic few-shot text classification (DFSTC) task. As shown in Figure 1, the DFSTC task begins with a few initial classes (e.g., Machine Learning). As new topics like Computer Vision emerge in later rounds, models must adapt to these changes and accurately classify new examples with minimal labeled data. Hence, how to develop methods to dynamically classify text with limited labeled data available remains a necessity and open problem.
Figure 1. An example of the DFSTC task with two rounds. In round 1, an example text is classified as Machine Learning. In round 2, an example text is classified as Computer Vision.
@@
Depending the technique for the DFSTC task, current models can be mainly categorized into two types, i.e., data augmentation-based models and RAG-based models. Firstly, data augmentation-based models (Meng et al., 2018; Meng et al., 2020; Xia et al., 2021) create additional training data by mixing the pairs of the few-shot labeled data and assigning mixed labels indicating the validity for these created data based on the labels of each data pairs (Xia et al., 2021). Also, several data augmentation approaches create extra semantic-related content based on the label names (Meng et al., 2018; Meng et al., 2020) to enrich the training corpus further. However, due to the limited labeled data, and the label names may not always be available, the generated text data can have very limited patterns. Consequently, the text classification models trained on these generated data may over-fit on limited text data and are not generalizable (Liu et al., 2023; Liu et al., 2024c; Lei et al., 2023).
Figure 2. An overview of GraphRAG (Edge et al., 2024). It first extracts and indexes entities and relations as graph nodes and edges. Then it performs Top-K similarity-based retrieval for query-related subgraphs. Finally, it classifies texts based on the subgraph summary and the query text.
@@
Recently, large language models (LLMs) (Sun et al., 2021; Jiang et al., 2023; Touvron et al., 2023; Achiam et al., 2023; Yang et al., 2024), pre-trained on extensive corpora, have achieved significant success due to their superior and comprehensive text understanding abilities. However, researchers (Gao et al., 2023; Peng et al., 2024) discover that LLMs can struggle to understand their inputs due to a lack of task-specific knowledge. Therefore, the Long Context RAG models (Chen et al., 2023; Sarthi et al., 2024) were proposed to combine the predicted text, candidate labels, and the retrieved side information (e.g., descriptions of text and labels, or documents) as LLM inputs, enable LLMs to understand the predicted text and labels more comprehensively with the help of side information from external sources. However, the incorporation of side information can further increase the input size and noise (Chen et al., 2023), which impedes the efficiency and effectiveness of LLMs (Sarthi et al., 2024).
As a result, Compression-based RAG approaches (Edge et al., 2024; Gutiérrez et al., 2024; Guo et al., 2024) are proposed to compress the length of context and extract the key information. Specifically, LongLLMLingua (Xia et al., 2021) and CompaAct (Yoon et al., 2024) propose filtering uninformative texts based on the query. However, these approaches may result in information loss and degrade overall performance. Then, to address this, graph RAG approaches (Edge et al., 2024; Guo et al., 2024; Gutiérrez et al., 2024; Peng et al., 2024; Zhang et al., 2025; Han et al., 2024a), such as GraphRAG (Edge et al., 2024) and LightRAG (Guo et al., 2024), propose to to index the unstructured texts into the structured graphs. For example, in Figure 2, the basic idea of graph RAG considers text entities as graph nodes, and text sentences describing relations between these entity nodes as graph edges, later, it retrieves subgraphs base on the query text, and generates short summaries of these subgraphs as the side information for LLMs, reducing the input length for LLMs. However, existing Compression RAG approaches (Edge et al., 2024; Guo et al., 2024; Gutiérrez et al., 2024) still have three issues in the DFSTC task.
  1. (1)
    Uniform-indexing issue: Some of these approaches construct graphs by indexing extracted text chunks uniformly. They do not consider the varying importance and extraction confidence of each text chunk, which may provide incorrect and unreliable context for LLMs.
  2. (2)
    Threshold-dependent issue: These approaches select relevant information for each input text based on a globally predefined threshold. However, the optimal retrieval threshold can vary across different text samples; hence, relying on a globally predefined threshold can be suboptimal for the entire dataset.
  3. (3)
    Narrow-source issue: These approaches only retrieve side information from the each few-shot training text and ignore the important information among different the query texts, consequently limiting their retrieve accuracy and comprehensiveness.
To address these issues, we propose a novel hyperparameter-free Graph-based Online Retrieval Augmented Generation framework for dynamic few-shot text classification, called GORAG. In general, GORAG constructs and maintains an adaptive online weighted graph by extracting side information from all target texts, and tailoring the graph retrieval for each input.
@@
Firstly, GORAG constructs a weighted graph with keywords extracted from labeled texts as keyword nodes, and text labels as label nodes. The edge of this graph represents the relationship between keywords and labels. Secondly, GORAG employs an edge-weighting mechanism to assign different weights to edges created by the last step. An example of the constructed weighted graph is shown in Figure 3. The edge weights represent the keywords’ importance and relevance to the respective text’s label. Thirdly, GORAG constructs a minimum-cost spanning tree on the constructed weighted graph based on keywords of the query text, and then retrieves the label nodes within the spanning tree as candidate labels. Since the generated spanning tree is determined solely by the constructed graph and the keywords of the query text, GORAG achieves adaptive retrieval without relying on any human-defined retrieval thresholds, making the retrieval more precise regarding each query text. Lastly, GORAG applies an online indexing mechanism, which can enrich the weighted graph with the keywords from query texts, enhancing its few-shot learning performance by providing a more comprehensive retrieve source.
We summarize our novel contributions as follows.
  • We present a hyperparameter-free online RAG framework for DFSTC tasks, namely GORAG. The GORAG framework consists of three steps: graph construction, candidate label retrieval, and text classification, aim at addressing the uniform-index, threshold-dependent and narrow-source issues.
@@
  • We develop a novel graph edge weighting mechanism based on the text keywords’ importance within the text corpus, which enables our approach to effectively model the relevance between keywords and labels, thereby helping the more accurate retrieval.
@@
  • To avoid any human-selected thresholds, we formulate the candidate label retrieval problem which is akin to the NP-hard Steiner Tree problem. To solve this problem, provide an efficient and effective solution to generate candidate labels.
@@ 2. Preliminary and Related Works
We first introduce the preliminaries of dynamic few-shot text classification in Section 2.1 and then discuss the related works for DFSTC task in Section 2.2. The important notations used in this paper are listed in Table 6 in Appendix A.
2.1. Dynamic Few-shot Text Classification Task
@@
Text classification (Kowsari et al., 2019; Minaee et al., 2021; Li et al., 2022) is a key task in real-world application that involves assigning predefined labels $y\in\mathcal{Y}$ to text $t=(w_{1},\cdots,w_{|t|})$ based on its words $w_{i}$. Traditional approaches require fine-tuning with large labeled datasets (Devlin, 2018; Liu, 2019), which may not always be available. Recently, few-shot learning addresses this limitation by enabling models to classify text with only a small number of labeled examples per class (Meng et al., 2018; Meng et al., 2020). On the other hand, dynamic classification (Parisi et al., 2019; Hanmo et al., 2024; Wang et al., 2024d) introduces an additional challenge where new classes are introduced over multiple rounds, requiring the model to adapt to new classes while retaining knowledge of previously seen ones. Combining these aspects, Dynamic Few-Shot Text Classification (DFSTC) task allows the model to handle evolving classification tasks with minimal labeled data, and better suited for real world scnerio (Xia et al., 2021).
@@
In DFSTC task, the model is provided with multiple rounds of new class updates. Specifically, In each round $r$, a new set of classes $\mathcal{Y}_{new}^{r}$ is introduced, and the labelled dataset for $\mathcal{Y}_{new}^{r}$ is denoted as $\mathcal{D}_{new}^{r}=\cup_{y_{i}\in\mathcal{Y}_{new}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$, where per class $y_{i}\in\mathcal{Y}_{new}^{r}$ only has $k$ labeled examples $\{t_{j},y_{i}\}_{j=1}^{k}$. Also, we denote candidate cumulative labels from the first round to the $r$-th round as $\mathcal{Y}^{r}=\bigcup_{i=1}^{r}\mathcal{Y}_{new}^{i}$. Formally, at the $r$-th round, given the candidate labels $\mathcal{Y}^{r}$ and all labeled data $\mathcal{D}^{r}=\cup_{i=1}^{r}\mathcal{D}_{new}^{i}$, the target of DFSTC task is to learn a function $f^{r}_{\theta}$, which can learn scores for all target labels $\mathbf{s}^{r}={f}^{r}_{\theta}(t_{u},\mathcal{Y}^{r})\in[0,1]^{|\mathcal{Y}^{r}|}$ for the unseen text $t_{u}$. Then, we can get the predicted label $y^{r*}\in\mathcal{Y}^{r}$ for the unseen text $t_{u}$ as follows.
(1) $y^{r*}=\arg\max_{y^{r}\in\mathcal{Y}^{r}}({f}_{\theta}^{r}(t_{u},\mathcal{Y}^{r})[y^{r}])$
DFSTC is valuable for its ability to learn from limited labeled data, adapt to evolving class distributions, and address real-world scenarios where classes and data evolve over time.
2.2. Dynamic Few-shot Text Classification Model
Current DFSTC models can be broadly categorized into NN-based models and RAG-based models, RAG-based models can be further categorized into Long Context RAG models, and Compression-based RAG models.
2.2.1. NN-based models